OpenStudy (anonymous):

Compute $\dfrac{(2/3)^3}{(2/9)^2}$.

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle \frac{\left(\dfrac{2}{3}\right)^3}{\left(\dfrac{2}{9}\right)^3} }\) like this?

OpenStudy (solomonzelman):

Here the $ don't work, you have to use `\( text/tex/something \)`

OpenStudy (solomonzelman):

Most general example of this: \(\large \color{#000000 }{ \displaystyle \frac{\left(\dfrac{a}{b}\right)^k}{\left(\dfrac{a}{b^{2}}\right)^k} }\) \(\large \color{#000000 }{ \displaystyle \frac{\quad \dfrac{a^k}{b^k}\quad }{\dfrac{a^k}{(b^{2})^k}} }\) \(\large \color{#000000 }{ \displaystyle\dfrac{a^k}{b^k}\div \dfrac{a^k}{(b^{2})^k}}\) \(\large \color{#000000 }{ \displaystyle\dfrac{a^k}{b^k}\times \dfrac{(b^{2})^k}{a^k}}\) \(\large \color{#000000 }{ \displaystyle\dfrac{(b^{2})^k}{b^k}}\) \(\large \color{#000000 }{ \displaystyle\dfrac{b^{2k}}{b^k}=b^{2k-k}=b^k}\)

OpenStudy (anonymous):

like that can u answer it

OpenStudy (solomonzelman):

I just gave you an abstract example...

OpenStudy (solomonzelman):

(A, b and k are some non-zero real numbers- just like the numbers in your problems)